# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ I.3 Over-specified Setting
$B_{n}$ $=\sum_{i=1}^{K}\sum_{|\gamma|=1}\frac{\exp(\beta_{i}^{n})}{\gamma!}(\Delta W_{i}^{n})^{\gamma}\cdot\dfrac{\partial^{|\gamma|}H}{\partial W^{\gamma}}(Y|X;W_{i}^{*})+R_{2}(X,Y)$
$=\sum_{i=1}^{K}\sum_{|\gamma|=1}\frac{\exp(\beta_{i}^{n})}{\gamma!}(\Delta W_{i}^{n})^{\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)+R_{2}(X,Y),$ (18)
@@
$Q_{n}$ $=\sum_{i=1}^{K}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}S^{n}_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{i}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})$
$+\sum_{i=1}^{K}\sum_{|\gamma|=0}^{1}T^{n}_{i,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)+R_{1}(X,Y)+R_{2}(X,Y),$ (19)
in which we respectively define for each $i\in[K]$ that
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\Big{|}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{|}=\sum_{i=1}^{K}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$ (20)
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\|\Delta W_{i}^{n}\|_{1}=\sum_{i=1}^{K}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\|\Delta a_{i}^{n}\|_{1}=\sum_{i=1}^{K}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\|\Delta b_{i}^{n}\|_{1}=\sum_{i=1}^{K}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\|\Delta\nu_{i}^{n}\|_{1}=\sum_{i=1}^{K}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{i}^{n}\|_{1}+\|\Delta a_{i}^{n}\|_{1}+|\Delta b_{i}^{n}|+|\Delta\nu_{i}^{n}|\Big{]}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{K}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{i}^{n}\|+\|\Delta a_{i}^{n}\|+|\Delta b_{i}^{n}|+|\Delta\nu_{i}^{n}|\Big{]}\to 0.$ (21)
@@
$\sum_{i=1}^{K}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\xi_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{i}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})$
$\hskip 199.16928pt+\sum_{i=1}^{K}\sum_{|\gamma|=0}^{1}\kappa_{i,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)=0,$ (22)
@@
$\sum_{i=1}^{K}\sum_{|\alpha_{1}|=0}^{1}\left[\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\xi_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\frac{\partial^{\alpha_{2}+\eta}f}{\partial h_{1}^{\alpha_{2}+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})+\kappa_{i,\alpha_{1}}p_{G_{*}}(Y|X)\right]$
@@
for any $i\in[K]$, $0\leq|\alpha_{1}|\leq 1$ and for almost surely $(X,Y)$. Moreover, since $(a^{*}_{1},b^{*}_{1},\nu^{*}_{1}),\ldots,(a^{*}_{K},b^{*}_{K},\nu^{*}_{K})$ have pair-wise distinct values, those of $((a^{*}_{1})^{\top}X+b^{*}_{1},\nu^{*}_{1}),\ldots,((a^{*}_{K})^{\top}X+b^{*}_{K},\nu^{*}_{K})$ are also pair-wise different. Therefore, the set
@@
is also linearly independent. Consequently, we obtain that $\xi_{i,\alpha_{1},\alpha_{2},\eta}=\kappa_{i,\gamma}=0$ for any $i\in[K]$, $0\leq|\alpha_{1}|+\alpha_{2}\leq 1$, $0\leq\eta\leq 2(1-|\alpha_{1}|-|\alpha_{2}|)$ and $0\leq|\gamma|\leq 1$, which contradicts the fact that at least one among those terms is different from zero.
@@
$\inf_{G\in\mathcal{G}_{k_{*}}(\Omega):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0.$
Proof of inequality B: Assume by contrary that the inequality B does not hold, then there exists a sequence of mixing measures $G^{\prime}_{n}\in\mathcal{G}_{k_{*}}(\Omega)$ such that $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$ and
@@
This result leads to $\mathbb{E}_{X}[V(p_{G^{\prime}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\to 0$ as $n\to\infty$. Recall that $\Omega$ is a compact set, therefore, we can replace the sequence $G^{\prime}_{n}$ by one of its subsequences that converges to a mixing measure $G^{\prime}\in\mathcal{G}_{k_{*}}(\Omega)$. Since $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, this result induces that $\mathcal{D}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}$.

### v4

@@ I.3 Over-specified Setting
$B_{n}$ $=\sum_{i=1}^{k_{*}}\sum_{|\gamma|=1}\frac{\exp(\beta_{i}^{n})}{\gamma!}(\Delta W_{i}^{n})^{\gamma}\cdot\dfrac{\partial^{|\gamma|}H}{\partial W^{\gamma}}(Y|X;W_{i}^{*})+R_{2}(X,Y)$
$=\sum_{i=1}^{k_{*}}\sum_{|\gamma|=1}\frac{\exp(\beta_{i}^{n})}{\gamma!}(\Delta W_{i}^{n})^{\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)+R_{2}(X,Y),$ (12)
@@
$Q_{n}$ $=\sum_{i=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}S^{n}_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{i}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})$
$+\sum_{i=1}^{k_{*}}\sum_{|\gamma|=0}^{1}T^{n}_{i,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)+R_{1}(X,Y)+R_{2}(X,Y),$ (13)
in which we respectively define for each $i\in[k_{*}]$ that
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\Big{|}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{|}=\sum_{i=1}^{k_{*}}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$ (14)
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\|\Delta W_{i}^{n}\|_{1}=\sum_{i=1}^{k_{*}}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\|\Delta a_{i}^{n}\|_{1}=\sum_{i=1}^{k_{*}}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\|\Delta b_{i}^{n}\|_{1}=\sum_{i=1}^{k_{*}}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\|\Delta\nu_{i}^{n}\|_{1}=\sum_{i=1}^{k_{*}}\frac{|S^{n}_{i,\alpha_{1},\alpha_{2},\eta}|}{\mathcal{D}_{1}(G_{n},G_{*})}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{i}^{n}\|_{1}+\|\Delta a_{i}^{n}\|_{1}+|\Delta b_{i}^{n}|+|\Delta\nu_{i}^{n}|\Big{]}\to 0.$
@@
$\frac{1}{\mathcal{D}_{1}(G_{n},G_{*})}\cdot\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{i}^{n}\|+\|\Delta a_{i}^{n}\|+|\Delta b_{i}^{n}|+|\Delta\nu_{i}^{n}|\Big{]}\to 0.$ (15)
@@
$\sum_{i=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\xi_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{i}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})$
$\hskip 199.16928pt+\sum_{i=1}^{k_{*}}\sum_{|\gamma|=0}^{1}\kappa_{i,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{i}^{*})p_{G_{n}}(Y|X)=0,$ (16)
@@
$\sum_{i=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{1}\left[\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\xi_{i,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\frac{\partial^{\alpha_{2}+\eta}f}{\partial h_{1}^{\alpha_{2}+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*})+\kappa_{i,\alpha_{1}}p_{G_{*}}(Y|X)\right]$
@@
for any $i\in[k_{*}]$, $0\leq|\alpha_{1}|\leq 1$ for almost surely $(X,Y)$. Moreover, since $(a^{*}_{1},b^{*}_{1},\nu^{*}_{1}),\ldots,(a^{*}_{k_{*}},b^{*}_{k_{*}},\nu^{*}_{k_{*}})$ have pair-wise distinct values, those of $((a^{*}_{1})^{\top}X+b^{*}_{1},\nu^{*}_{1}),\ldots,((a^{*}_{k_{*}})^{\top}X+b^{*}_{k_{*}},\nu^{*}_{k_{*}})$ are also pair-wise different. Therefore, the set
$\Big{\{}X^{\alpha_{2}}\frac{\partial^{\alpha_{2}+\eta}f}{\partial h_{1}^{\alpha_{2}+\eta}}(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*}),\ p_{G_{*}}(Y|X):$
@@
is also linearly independent. Consequently, we obtain that $\xi_{i,\alpha_{1},\alpha_{2},\eta}=\kappa_{i,\gamma}=0$ for any $i\in[k_{*}]$, $0\leq|\alpha_{1}|+\alpha_{2}\leq 1$, $0\leq\eta\leq 2(1-|\alpha_{1}|-|\alpha_{2}|)$ and $0\leq|\gamma|\leq 1$, which contradicts the fact that at least one among those terms is different from zero.
@@
$\inf_{G\in\mathcal{G}_{k_{*}}(\Theta):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0.$
Proof of inequality B: Assume by contrary that the inequality B does not hold, then there exists a sequence of mixing measures $G^{\prime}_{n}\in\mathcal{G}_{k_{*}}(\Theta)$ such that $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$ and
@@
This result leads to $\mathbb{E}_{X}[V(p_{G^{\prime}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\to 0$ as $n\to\infty$. Recall that $\Omega$ is a compact set, therefore, we can replace the sequence $G^{\prime}_{n}$ by one of its subsequences that converges to a mixing measure $G^{\prime}\in\mathcal{G}_{k_{*}}(\Theta)$. Since $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, this result induces that $\mathcal{D}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}$.
@@
K.2 Proof of Theorem 3.1
In this appendix, we employ results for M-estimators in van de Geer [2000] to establish the density estimation rate under the Laplace gating Gaussian mixture of experts (MoE).
[…]
Lemma K.1.
If $\Theta$ is a bounded set, then the following inequalities hold for any $20<\eta<1/$ 2:
  • (i)
    $\log N(\eta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})\lesssim\log(1/\eta)$;
  • (ii)
    $H_{B}(\eta,\mathcal{P}_{k}(\Theta),h)\lesssim\log(1/\eta)$.
Proof of Lemma K.1 is in Appendix K.2.2. Subsequently, we denote
$\widetilde{\mathcal{P}}_{k}(\Theta)$ :=$:=\{p_{(G+G_{*})/2}(Y|X):G\in\mathcal{G}_{k}(\Theta)\};$
$\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta)$ :=$:=\{p^{1/2}_{(G+G_{*})/2}(Y|X):G\in\mathcal{G}_{k}(\Theta)\}.$
In addition, for each $\delta>0$, we define a Hellinger ball centered around the conditional density function $p_{G_{*}}(Y|X)$ and intersected with the set $\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta)$ as
$\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,\delta):=\{p^{1/2}\in\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta):h(p,p_{G_{*}})\leq\delta\}.$
To capture the size of the above Hellinger ball, van de Geer [2000] suggest using the following quantity:
$\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,\delta)):=\int_{\delta^{2}/2^{13}}^{\delta}H^{1/2}_{B}(t,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,t),\|\cdot\|_{2})\mathrm{d}t\vee\delta,$ (17)
where $t\vee\delta:=\max\{t,\delta\}$. Given those notations, let us recall a standard result for density estimation in van de Geer [2000].
Lemma K.2 (Theorem 7.4, van de Geer [2000]).
Take $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,\delta))$ such that $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Then, for some sequence $(\delta_{n})$ and universal constant $c$ which satisfy $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta)$, we obtain that
for any $\delta\geq\delta_{n}$
Proof of Lemma K.2 can be found in van de Geer [2000]. Now, we are ready to provide the proof for convergence rate of density estimation in Theorem J.1 in Appendix K.2.1.
K.2.1 Main Proof
It is worth noting that for any $t>0$, we have
$H_{B}(t,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,t),\|\cdot\|_{2})\leq H_{B}(t,\mathcal{P}_{k}(\Omega,t),h).$
Then, the integral in equation 17 is upper bounded as follows:
$\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,\delta))\leq\int_{\delta^{2}/2^{13}}^{\delta}H^{1/2}_{B}(t,\mathcal{P}_{k}(\Omega,t),h)\mathrm{d}t\vee\delta\lesssim\int_{\delta^{2}/2^{13}}^{\delta}\log(1/t)\mathrm{d}t\vee\delta,$ (18)
where the second inequality follows from part (ii) of Lemma K.1.
As a result, by choosing $\Psi(\delta)=\delta\cdot\sqrt{\log(1/\delta)}$, we can verify that $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, the inequality in equation 18 indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k}(\Theta,\delta))$. Next, let us consider a sequence $(\delta_{n})$ defined as $\delta_{n}:=\sqrt{\log(n)/n}$. This sequence can be validated to satisfy the condition $\sqrt{n}\delta_{n}^{2}\geq c\Psi(\delta)$ for some universal constant $c$. Therefore, by Lemma K.2, we reach the conclusion of Theorem J.1:
for some universal constant $C$ depending only on $\Theta$.
K.2.2 Proof of Lemma K.1
Part (i). In this part, we will derive the following upper bound for the covering number of metric space $(\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})$ for any $20<\eta<1/$ 2 given the bounded set $\Omega$:
$\log N(\eta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})\lesssim\log(1/\eta).$
To start with, we denote $\Omega:=\{(a,b,\nu)\in\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}_{+}:(\beta,W,a,b,\nu)\in\Omega\}$. As $\Theta$ is a bounded set, the set $\Omega$ is also bounded. Therefore, we can find an $\eta$-cover of $\Omega$, denoted by $\overline{\Omega}_{\eta}$. Additionally, we also define $\Delta:=\{(\beta,W)\in\mathbb{R}\times\mathbb{R}^{d}:(\beta,W,a,b,\nu)\in\Omega\}$, and $\overline{\Delta}_{\eta}$ be an $\eta$-cover of $\Delta$. Then, it can be validated that
$|\overline{\Omega}_{\eta}|\leq\mathcal{O}(\eta^{-(d+2)k}),\quad|\overline{\Delta}_{\eta}|\leq\mathcal{O}(\eta^{-(d+1)k}).$
Next, for each mixing measure $G=\sum_{i=1}^{k}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}\in\mathcal{G}_{k}(\Theta)$, we take into account two other mixing measures. The first measure is $G^{\prime}=\sum_{i=1}^{k}\exp(\beta_{i})\delta_{(W_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}$, where $(\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Omega}_{\eta}$ is the closest points to $(a_{i},b_{i},\nu_{i})$ in this set for all $i\in[k]$. The second one is $\overline{G}:=\sum_{i=1}^{k}\exp(\overline{\beta}_{i})\delta_{(\overline{W}_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}$ in which $(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta}$ for any $i\in[k]$. Next, let us define
$\mathcal{T}:=\{p_{\overline{G}}\in\mathcal{P}_{k}(\Theta):(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta},\ (\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Theta}_{\eta},\forall i\in[k]\},$
then it is obvious that $p_{\overline{G}}\in\mathcal{T}$. Now, we will show that $\mathcal{T}$ is an $\eta$-cover of metric space $(\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})$ with a note that it is not necessarily the smallest cover. Indeed, according to the triangle inequality, we have
$\|p_{G}-p_{\overline{G}}\|_{\infty}\leq\|p_{G}-p_{G^{\prime}}\|_{\infty}+\|p_{G^{\prime}}-p_{\overline{G}}\|_{\infty}.$ (19)
$\|p_{G}-p_{G^{\prime}}\|_{\infty}$ $\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\mathrm{softmax}(-\|W_{i}-X\|+\beta_{i})\cdot\Big{|}f(Y|a_{i}^{\top}X+b_{i},\nu_{i})-f(Y|\overline{a}_{i}^{\top}X+\overline{b}_{i},\overline{\nu}_{i})\Big{|}$
$\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\Big{|}f(Y|a_{i}^{\top}X+b_{i},\nu_{i})-f(Y|\overline{a}_{i}^{\top}X+\overline{b}_{i},\overline{\nu}_{i})\Big{|}$
$\lesssim\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\Big{(}\|a_{i}-\overline{a}_{i}\|+\|b_{i}-\overline{b}_{i}\|+\|\nu_{i}-\overline{\nu}_{i}\|\Big{)}$
$=\sum_{i=1}^{k}\Big{(}\|a_{i}-\overline{a}_{i}\|+\|b_{i}-\overline{b}_{i}\|+\|\nu_{i}-\overline{\nu}_{i}\|\Big{)}$
$\lesssim\eta.$ (20)
Subsequently, we bound the second term $\|p_{G^{\prime}}-p_{\overline{G}}\|_{\infty}$ as follows:
$\|p_{G^{\prime}}-p_{\overline{G}}\|_{\infty}$ $\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\Big{\{}\Big{|}\mathrm{softmax}(-\|W_{i}-X\|+\beta_{i})-\mathrm{softmax}(-\|\overline{W}_{i}-X\|+\overline{\beta}_{i})\Big{|}$
×$\hskip 170.71652pt\times\Big{|}f(Y|\overline{a}_{\tau_{i}}^{\top}X+\overline{b}_{\tau_{i}},\overline{\nu}_{\tau_{i}})\Big{|}\Big{\}}$
$\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\Big{|}\mathrm{softmax}(-\|W_{i}-X\|+\beta_{i})-\mathrm{softmax}(-\|\overline{W}_{i}-X\|+\overline{\beta}_{i})\Big{|}$
$\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}\Big{|}-\|W_{i}-X\|+\beta_{i}+\|W_{i}-X\|-\beta_{i}\Big{|}$
$\leq\sum_{i=1}^{k}\sup_{X\in\mathcal{X}}[\|W_{i}-\overline{W}_{i}\|+|\beta_{i}-\overline{\beta}_{i}|]$
$\lesssim\eta,$ (21)
It follows from the results in equation 19, equation 20 and equation 21 that $\|p_{G}-p_{\overline{G}}\|_{\infty}\lesssim\eta$. This result indicates that $\mathcal{T}$ is an $\eta$-cover of the metric space $(\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})$. As a consequence, we obtain that
$N(\eta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})\lesssim|\overline{\Delta}_{\eta}|\times|\overline{\Omega}_{\eta}|\leq\mathcal{O}(1/\eta^{(2d+3)k}),$
which leads to the conclusion of this part: $\log N(\eta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})\lesssim\log(1/\eta)$.
Part (ii). In this part, we provide an upper bound for the bracketing entropy of $\mathcal{P}_{k}(\Theta)$ under the Hellinger distance $h$:
$H_{B}(\eta,\mathcal{P}_{k}(\Theta),h)\lesssim\log(1/\eta).$
Since $\Theta$ and $\mathcal{X}$ are bounded sets, there exist positive constants $\gamma,\ell,u$ such that $-\gamma\leq a^{\top}X+b\leq\gamma$ and $\ell\leq\nu\leq u$. Let us define
Next, let $\zeta\leq\eta$ which will be chosen later and $\{p_{1},\ldots,p_{N}\}$ be an $\zeta$-cover of metric space $(\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})$ with the covering number $N:=N(\zeta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})$. Additionally, we also consider brackets of the form $[\Psi_{i}^{L}(Y|X),\Psi_{i}^{U}(Y|X)]$ where
$\Psi_{i}^{L}(Y|X)$ $:=\max\{p_{i}(Y|X)-\zeta,0\}$
$\Psi_{i}^{U}(Y|X)$ $:=\max\{p_{i}(Y|X)+\zeta,B(Y|X)\}.$
Then, we can check that $\mathcal{P}_{k}(\Theta)\subseteq\bigcup_{i=1}^{N}[\Psi_{i}^{L}(Y|X),\Psi_{i}^{U}(Y|X)]$ and $\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)\leq\min\{2\zeta,B(Y|X)\}$.
$\|\Psi_{i}^{U}-\Psi_{i}^{L}\|_{1}$ $=\int_{|Y|<2\gamma}[\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)]~{}\mathrm{d}(X,Y)+\int_{|Y|\geq 2\gamma}[\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)]~{}\mathrm{d}(X,Y)$
$\leq S\zeta+\exp\Big{(}-\frac{S^{2}}{2u}\Big{)}\leq S^{\prime}\zeta,$
$H_{B}(S^{\prime}\zeta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{1})\leq\log N(\zeta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{\infty})\leq\log(1/\zeta).$
By setting $\zeta=\eta/S^{\prime}$, we obtain that $H_{B}(\eta,\mathcal{P}_{k}(\Theta),\|\cdot\|_{1})\lesssim\log(1/\eta)$. Finally, due to the inequality $h^{2}\leq\|\cdot\|_{1}$, we reach the conclusion of this part:
$H_{B}(\eta,\mathcal{P}_{k}(\Theta),h)\lesssim\log(1/\eta).$
Hence, the proof is completed.